The perimeter of a rectangular room is 37 3/5 ft. The width is 6 4/5 feet. Find the length.
I know the formula is p=2l + 2w but the fractions and information are confusing

Answer by gonzo(654)

P = 2*L + 2*W
37+(3/5) = 2*L + 2*(6+(4/5))
making the improper fraction into proper fraction gets
188/5 = 2*L + (2*(34/5))
this becomes
188/5 = 2*L + (2*34)/5
multiplying both sides of the equation by 5 to remove the denominators and
the equation becomes
188 = 2*5*L + (2*34)
which becomes
188 = 10*L + 68
subtracting 68 from both sides and equation becomes
188-68 = 10*L
which becomes
120 = 10*L
which becomes L = 12
-----
substituting in the original equation gets
37 + (3/5) = (2*12) + 2*(6 + 4/5)
converting to proper fraction and equation becomes
188/5 = 2*12 + 2*(34/5)
multiplying both sides by 5 to remove the denominators and equation becomes
188 = 2*5*12 + 2*34
which becomes
188 = 120 + 68
which becomes
188 = 188
so answer is good.